OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4).

OpenStudy (anonymous):

well just explain how I'm supposed to do this

OpenStudy (anonymous):

solve the equation for y. 4x-3y=-12 -4x -4x subtract 4y from both sides -3y =-4x-12 ___ ________ -3 -3 Divide negative 3 from both sides y = 4/3x +4 the solve of this is 4/3 do parallel slope is 4/3 now I have slope of my need line and two points slope: 4/3 and points (-3.4) can you do from here

OpenStudy (anonymous):

choices are y = 4/3x + 8 y = 4/3x + 3 y = −3/4x + 8 y = −3/4x + 3